# Supervised Graph Contrastive Learning for Gene Regulatory Networks

### v4

@@ 5 Experiments
Table 1: Description of the downstream tasks
Task Task Type Metrics
Node-Level Task
[Rel.]: Cancer relation Classification (binary) Accuracy
Graph-Level Task
This chapter verifies the effectiveness of the proposed method using actual GRNs from cancer patients and biologically augmented GRNs based on gene knockdown experiments.
5.1 Benchmark of Gene Regulatory Networks
Evaluation Protocol:
We evaluated the proposed method through the following procedure. First, we constructed patient-specific GRNs from cancer patient gene expression data and teacher GRNs from gene knockdown experiment data. Then, pre-training was performed on the proposed method using both the patient-specific and teacher GRNs. Subsequently, the performance of the downstream tasks, such as classification accuracy and regression performance, was evaluated using the pre-trained models and compared against comparative methods.
We compared the proposed method with the following five comparative models:
w/o-pretrain
: Directly performs classification or regression for downstream tasks.
GAE
(Kipf & Welling, 2016): Reconstruction based graph representation learning method.
GraphCL
(You et al., 2021): Graph contrastive learning using positive pairs between graphs.
GRACE
(Zhu et al., 2020): Node-level graph contrastive learning method.
SGRL
(He et al., 2024): State-of-the-art augmentation-free GCL.
Table 2: Fine-tuning results of node-level downstream tasks
@@ 5.1 Benchmark of Gene Regulatory Networks
Table 3: Fine-tuning results of graph-level downstream tasks
Task w/o-pretrain GAE GraphCL GRACE SGRL SupGCL
Hazard
@@
To evaluate the performance of SupGCL, we conducted benchmark evaluations using real-world datasets. For constructing patient-specific GRNs, we used cancer cell sample data from The Cancer Genome Atlas (TCGA). For constructing teacher GRNs, we used gene knockdown experiment data from cancer cell lines in the Library of Integrated Network-based Cellular Signatures (LINCS). The TCGA dataset (Weinstein et al., 2013) and the LINCS dataset  (Subramanian et al., 2017) are both large-scale and widely-used public platforms providing gene expression data from cancer patients and cell lines, respectively.
Experiments were conducted for three cancer types across both datasets: breast cancer, lung cancer, and colorectal cancer. Furthermore, the set of genes constituting each network was restricted to the 975 genes common to the TCGA gene set and the 978 LINCS landmark genes. The number of patient samples for each cancer type was N=1092 (breast), 1011 (lung), and 288 (colorectal), and the total number of knockdown experiments was 8793, 11843, and 15926, respectively. The number of unique knockdown target genes / total common genes was 768/975, 948/975, and 948/975.
The TCGA dataset also includes survival status and disease subtype labels associated with each gene expression profile. Additionally, each gene was annotated with multi-labels based on Gene Ontology  (Ashburner et al., 2000) — Biological Process (metabolism, signaling, cell organization; 3 classes), and Cellular Component (nucleus, mitochondria, ER, membrane; 4 classes). We also used the OncoKB  (Chakravarty et al., 2017) cancer-related gene list to assign binary relevance labels. These labels were used for downstream tasks. Details are provided in Appendix C.
@@
Figure 3: t-SNE visualization of pre-trained graph-level embeddings on breast cancer GRNs. Each point represents the readout feature of an individual patient’s network. NMI and ARI scores indicate quantitative clustering metrics of the embeddings.
5.2 Result 1: Evaluation by Downstream Task
@@
Graph-level tasks (hazard prediction, subtype classification) used survival and subtype labels from TCGA. Note that subtype classification was conducted only for breast cancer. Node-level tasks (Biological Process - BP., Cellular Component - CC., and cancer relevance - Rel.) used gene-level annotations from Gene Ontology and OncoKB. Details of these downstream tasks are summarized in Table 1.
@@
Each downstream task — hazard prediction, subtype classification, BP, and CC classification — was evaluated using 10-fold cross-validation. For cancer gene classification, due to label imbalance, we performed undersampling over 10 random seeds. Results are reported as mean ± standard deviation.
For all methods including the SupGCL and conventional methods, we used the same 5-layer Graph Transformer architecture (Shi et al., 2021). Hyperparameters for pre-training were tuned with Optuna (Akiba et al., 2019), and the model was optimized using the AdamW optimizer (Loshchilov & Hutter, 2019). All training runs were performed on a single NVIDIA H100 SXM5 GPU. Additional experimental details can be found in Appendix E.
Tables 2 and 3 show the results for node-level and graph-level tasks, respectively. The best performance is indicated in bold, and the second-best is underlined. Although SupGCL did not achieve statistically significant superiority in every single task, it consistently outperformed other pre-training methods across all datasets and tasks.
For node-level tasks, many existing methods did not show much improvement over without-pretrain, whereas SupGCL consistently demonstrated strong performance.
In graph-level tasks like hazard and subtype prediction, while some existing methods showed marginal improvement over without-pretrain, SupGCL achieved significantly higher performance.
In addition, we evaluated the generalization performance across different cancer types. While the pre-training was robust even when the cancer types of the teacher and patient GRNs differed, the fine-tuning was not (see Appendix G for details).
5.3 Result 2: Effectiveness of Supervised Learning
@@
In this section, to verify the effectiveness of the reference model $p_{\phi}$, we conducted additional experiments by varying the augmentation-level temperature parameter $\tau_{\rm a}$, which controls its degree of influence. This experiment also serves to empirically validate the theoretical relationship (Corollary 1) that as $\tau_{\rm a}$ increases, the SupGCL loss function asymptotically approaches that of unsupervised node-level GCL like GRACE. The experimental results for breast cancer in Table 4 show that SupGCL outperforms GRACE in all settings, demonstrating the effectiveness of incorporating biologically plausible information. Furthermore, as $\tau_{\rm a}$ increases and the influence of supervised information weakens, the performance of SupGCL tends to approach that of the unsupervised method GRACE, which supports Corollary 1. It should be noted that due to the different definitions of the denominator in their loss functions, this experiment and GRACE are not strictly identical.
$\tau_{\rm a}$ Hazard Subtype BP. $C$. Rel.
0.10 $\mathbf{0.670\pm 0.078}$ $0.837\pm 0.029$ $\mathbf{0.262\pm 0.035}$ $0.289\pm 0.049$ $0.586\pm 0.047$
0.25 (default) $0.650\pm 0.059$ $\mathbf{0.847\pm 0.036}$ $0.243\pm 0.052$ $\mathbf{0.291\pm 0.026}$ $0.600\pm 0.057$
0.50 $0.648\pm 0.053$ $0.846\pm 0.032$ $0.261\pm 0.034$ $0.284\pm 0.061$ $\mathbf{0.606\pm 0.039}$
1.00 $0.640\pm 0.056$ $0.835\pm 0.031$ $0.244\pm 0.042$ $0.280\pm 0.032$ $0.596\pm 0.048$
2.00 $0.656\pm 0.060$ $0.842\pm 0.031$ $0.237\pm 0.024$ $0.277\pm 0.058$ $0.590\pm 0.044$
Reference: GRACE $0.642\pm 0.064$ $0.841\pm 0.026$ $0.230\pm 0.051$ $0.236\pm 0.026$ $0.575\pm 0.035$
5.4 Result 3: Latent Space Analysis
We visualized the graph-level embedding spaces from the breast cancer dataset (Figure 3; additional visualizations including node-level embeddings are in Appendix F). The embeddings, colored by disease subtype, show that GAE and GraphCL fail to separate subtypes, while GRACE and SGRL exhibit moderate separation. SupGCL achieves the clearest separation, indicating a superior ability to learn subtype-specific representations. This qualitative observation was confirmed by quantitative clustering metrics. SupGCL consistently outperformed other methods on both the Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), affirming that its embeddings more effectively capture subtype-specific structure.
@@ 6 Conclusion
In this study, we proposed a supervised graph contrastive learning method, SupGCL, for representation learning of gene regulatory networks (GRNs), which incorporates real-world genetic perturbation data as supervision during training. We formulated GCL with supervision-guided augmentation selection within a unified probabilistic framework, and theoretically demonstrated that conventional GCL methods are special cases of our proposed formulation. Through benchmark evaluations using downstream tasks based on both node-level and graph-level embeddings of GRNs from cancer patients, SupGCL consistently outperformed existing GCL methods.
A key limitation of our study is that SupGCL, when pre-trained on a specific cancer type, fails to improve performance on downstream tasks for other cancer types. As future work, we plan to expand the target cancer types and develop a large-scale, general-purpose SupGCL model that can operate across multiple cancer types.
@@ Acknowledgments
This work was supported by JST Moonshot R&D Grant Number JPMJMS2021, JPMJMS2024, and JSPS KAKENHI Grant Number 25K00148. This paper is also based on results obtained from a project, JPNP14004, commissioned by the New Energy and Industrial Technology Development Organization (NEDO) This work used computational resources of the supercomputer Fugaku provided by RIKEN through the HPCI System Research Project (Project IDs: hp150272, ra000018) and TSUBAME4.0 supercomputer at Institute of Science Tokyo.
@@
Appendix Contents
Appendix A Detail Proof of Theorem 1
@@ Appendix B Algorithm
Our algorithm follows a standard training loop, consisting of the calculation of $\text{Loss}_{\rm SupGCL}$ and the optimization of $f_{\phi}$ using AdamW. Furthermore, to reduce computational costs, we employ sampling-based estimation of the normalization constant for the calculation of softmax functions $p_{\phi}(b|a)$ and $q_{\phi}(b|a)$, and use importance sampling for the calculation of $\text{Loss}_{\rm node}^{a,b}$ and $\text{Loss}_{\rm Aug}$.
@@
1:Graph Neural Net $f_{\phi}$, all patient GRNs $\mathcal{G}_{\rm all}=\{\mathcal{G}^{(i)}\}_{i=1}^{N}$, all teacher GRNs $\mathcal{H}_{\rm all}=\{\mathcal{H}_{a}^{(i)}\}_{i,\in\mathcal{I}_{a},a\in\mathcal{K}}$
2:for $\mathcal{G}\subset\mathcal{G}_{\rm all}$ do
3:  $a,b\sim\mathrm{U}_{\mathcal{K}}$
[…]
6:  ${\bm{Z}}^{a},{\bm{Z}}^{b}\leftarrow f_{\phi}(\mathcal{G}_{a}),~f_{\phi}(\mathcal{G}_{b})$ $\triangleright$ Embedding target GRNs
7:  ${\bm{Y}}^{a},{\bm{Y}}^{b}\leftarrow f_{\phi}(\mathcal{H}_{a}),~f_{\phi}(\mathcal{H}_{b})$ $\triangleright$ Embedding teacher GRNs
8:  $q_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{Z}}^{a},{\bm{Z}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ $\triangleright$ Calculate augmentation-level target model
9:  $p_{\phi}(b|a)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{Y}}^{a},{\bm{Y}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ $\triangleright$ Calculate augmentation-level teacher model
10:  $\text{Loss}_{\rm Aug}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\big(\log p_{\phi}(b|a)-\log q_{\phi}(b|a)\big)$ $\triangleright$ Importance sampling of $\text{Loss}_{\rm Aug}^{a,b}$
11:  $q_{\phi}(j|i,a,b)\leftarrow\text{softmax}\left(\Big[\frac{\langle{\bm{z}}^{a}_{i},{\bm{z}}^{b}_{*}\rangle}{\tau_{n}}\Big]\right)[j]$ $\triangleright$ Calculate node-level target model
12:  $\text{Loss}_{\rm node}^{a,b}\leftarrow\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\log q_{\phi}(i|i,a,b)$
13:  $\text{Loss}_{\rm SupGCL}\leftarrow|\mathcal{K}|p_{\phi}(b|a)\text{Loss}_{\rm node}^{a,b}+\text{Loss}_{\rm Aug}$ $\triangleright$ Importance sampling of $\text{Loss}_{\rm node}^{a,b}$
14:  Update $f_{\phi}$ using $\text{Loss}_{\rm SupGCL}$ and AdamW optimizer
15:end for
16:return $\text{Trained Graph NN:}~f_{\phi}$

### v5

@@ 5 Experiments
Figure 3: t-SNE visualization of pre-trained graph-level embeddings on breast cancer GRNs. Each point represents the readout feature of an individual patient’s network. NMI and ARI scores indicate quantitative clustering metrics of the embeddings across 5 experimental runs.
Table 1: Fine-tuning results on downstream tasks. BP, $C$, and Rel are node-level tasks, while Hazard and Subtype are graph-level tasks. Breast, Lung, and Colorectal denote cancer types.
@@
Hazard [C-index $\uparrow$]
@@
5.1 Benchmark of Gene Regulatory Networks
Evaluation Protocol:
We evaluated the proposed method through the following procedure. First, we constructed patient-specific GRNs from cancer patient gene expression data and teacher GRNs from gene knockdown experiment data.
Second, we pre-trained the proposed model on patient-specific and teacher GRNs, and evaluated the embedding space using NMI and ARI with respect to disease subtype clustering. Finally, we evaluated 13 fine-tuning downstream task performance (e.g., classification accuracy and regression performance) using the pre-trained models. The compared five baselines listed below.
w/o-pretrain: Directly performs classification or regression for downstream tasks.
GAE(Kipf & Welling, 2016): Reconstruction based graph representation learning method.
GraphCL(You et al., 2021): Graph contrastive learning method using positive pairs between graphs.
GRACE(Zhu et al., 2020): Node-level graph contrastive learning method.
SGRL (He et al., 2024): State-of-the-art (SOTA) augmentation-free GCL method.
These five models consist of representative methods for four types of graph representation learning model, together with a SOTA GCL model. Other comparison results (augmentation-free GCLs, augmentation-adjustment GCLs, and biological foundation models) are presented in Appendix K.
@@ 5.1 Benchmark of Gene Regulatory Networks
We benchmarked on real-world datasets, constructing patient-specific GRNs from cancer samples in The Cancer Genome Atlas (TCGA)  (Weinstein et al., 2013) and teacher GRNs from gene knockdown experiments in cancer cell lines from the Library of Integrated Network-based Cellular Signatures (LINCS)  (Subramanian et al., 2017a). The TCGA and the LINCS are both large-scale and widely-used public platforms providing gene expression dataset from cancer patients and cell lines, respectively.
We considered three cancer types available in both resources: breast, lung, and colorectal cancer. To ensure a consistent node set across TCGA- and LINCS-derived networks, we restricted all GRNs to the intersection of genes measured in TCGA and the LINCS L1000 landmark set, resulting in 975 genes. TCGA included N=1092 (breast), 1011 (lung), and 288 (colorectal) patient samples. For LINCS, the total numbers of knockdown experiments were 8793 (breast), 11843 (lung), and 15926 (colorectal); among the 975 common genes, the numbers of unique knockdown targets were 768, 948, and 948, respectively.
Beyond expression profiles, we used additional annotations for downstream evaluation. TCGA provides survival status and disease subtype labels for each patient sample.
At the gene level, we assigned multi-label Gene Ontology (GO) annotations—Biological Process (BP; metabolism, signaling, cell organization; 3 classes) and Cellular Component (CC; nucleus, mitochondria, ER, membrane; 4 classes)—as well as binary cancer-relevance labels based on the OncoKB cancer-related gene list (Rel). These labels were used for downstream tasks. (see Appendices C and E in detail.)
@@
Model Setting:
For all methods including the SupGCL and comparative models, we used the same 5-layer Graph Transformer architecture (Shi et al., 2021). Hyperparameters for pre-training were tuned with Optuna (Akiba et al., 2019), and the model was optimized using the AdamW optimizer (Loshchilov & Hutter, 2019). All training runs were performed on a single NVIDIA H100 SXM5 GPU.
Additional experimental details are provided in Appendix E, and empirical runtime and computational complexity analyses are reported in Appendix F.
[…]
0.10 $\mathbf{0.262\pm 0.035}$ $0.289\pm 0.049$ $0.586\pm 0.047$ $\mathbf{0.670\pm 0.078}$ $0.837\pm 0.029$
0.25 (default) $0.243\pm 0.052$ $\mathbf{0.291\pm 0.026}$ $0.600\pm 0.057$ $0.650\pm 0.059$ $\mathbf{0.847\pm 0.036}$
0.50 $0.261\pm 0.034$ $0.284\pm 0.061$ $\mathbf{0.606\pm 0.039}$ $0.648\pm 0.053$ $0.846\pm 0.032$
1.00 $0.244\pm 0.042$ $0.280\pm 0.032$ $0.596\pm 0.048$ $0.640\pm 0.056$ $0.835\pm 0.031$
2.00 $0.237\pm 0.024$ $0.277\pm 0.058$ $0.590\pm 0.044$ $0.656\pm 0.060$ $0.842\pm 0.031$
Reference: GRACE $0.230\pm 0.051$ $0.236\pm 0.026$ $0.575\pm 0.035$ $0.642\pm 0.064$ $0.841\pm 0.026$
5.2 Result 1: Embedding Space Analysis
We visualized the graph-level embedding spaces from the breast cancer type (Figure 3). The embeddings, colored by disease subtype, show that GAE and GraphCL fail to separate subtypes, while GRACE and SGRL exhibit moderate separation. SupGCL significantly outperforming other methods on both the Normalized Mutual Information (NMI) and the Adjusted Rand Index (ARI), affirming that its embeddings capture subtype-specific structure. This advantage is statistically supported by quantitative metrics across 5 pre-training runs.
Furthermore, we evaluated the biological validity of the node-level embeddings performing enrichment analysis (Appendix  J). These analyses demonstrated that SupGCL captured the biological context more effectively than existing methods.
5.3 Result 2: Evaluation by Downstream Task
@@
Graph-level tasks (hazard prediction, subtype classification) used survival and subtype labels from TCGA. Note that subtype classification was conducted only for breast cancer.
Node-level tasks (BP., CC., and Rel.) used gene-level annotations from Gene Ontology and OncoKB.
@@
Each downstream task was evaluated using 10-fold cross-validation. For cancer gene classification, due to label imbalance, we performed undersampling over 10 random seeds. The results are reported as mean $\pm$ standard deviation.
Table 1 reports the results of node-level and graph-level downstream tasks, with the best and second-best performances indicated in bold and underlined, respectively. SupGCL achieves the highest performance across all datasets and tasks compared to other pre-training methods. This performance advantage extends to a broader range of methods, including augmentation-free and augmentation-adjustment GCL methods and biological foundation models, as detailed in Appendix K.
Note that for the node-level task, existing graph representation learning methods exhibit little difference compared to the results without pretraining. This indicates that graph representation learning methods without a teacher GRN fail to learn meaningful representations that capture gene-level characteristics relevant to this downstream task. In contrast, our proposed method leverages a teacher GRN, enabling the learning of gene representations associated with biological perturbations, which leads to improved performance.
In addition, we evaluated the generalization ability and robustness of SupGCL (see Appendices H and I, respectively). SupGCL demonstrated that pre-training remained effective even when the cancer types of the teacher and patient GRNs differed. Furthermore, our model exhibited strong robustness under limited availability of patient or teacher GRNs, as well as in the presence of noise in the estimated GRNs.
5.4 Result 3: Effectiveness of Supervised Learning
@@
In this section, to verify the effectiveness of the reference model $p_{\phi}$, we conducted additional experiments by varying the augmentation-level temperature parameter $\tau_{\rm a}$, which controls its degree of influence. This experiment also serves to empirically validate the theoretical relationship (Corollary 4.2) that as $\tau_{\rm a}$ increases, the SupGCL loss function asymptotically approaches that of unsupervised node-level GCL like GRACE. The experimental results for breast cancer in Table 2 show that SupGCL outperforms GRACE in all settings, demonstrating the effectiveness of incorporating biologically plausible information. Furthermore, as $\tau_{\rm a}$ increases and the influence of supervised information weakens, the performance of SupGCL tends to approach that of the unsupervised method GRACE, which supports Corollary 4.2. It should be noted that due to the different definitions of the denominator in their loss functions, this experiment and GRACE are not strictly identical.
@@ 6 Conclusion
In this study, we proposed SupGCL, a supervised graph contrastive learning method for Gene Regulatory Networks (GRNs) that incorporates biological perturbations from real-world gene knockdown experiments as supervision for graph augmentation. This proposed method formulates supervised graph augmentation within a unified probabilistic framework and theoretically derived that existing node-level GCL methods can be derived as special cases of our formulation.
In our empirical evaluation, an analysis of the learned embedding space first confirmed that SupGCL significantly improves clustering metrics (NMI/ARI) for cancer subtypes, capturing latent structures with high biological plausibility. Furthermore, evaluating the pre-trained model on downstream tasks with fine-tuning, SupGCL consistently outperformed state-of-the-art baselines across 13 tasks, including gene function classification and patient survival analysis.
A limitation of our study is the challenge of cross-domain generalization caused by distributional shifts across different cancer types. However, preliminary results presented in the Appendix H suggest that pan-cancer learning—integrating multiple cancer types—is effective, enriching representations without compromising performance. Future work will focus on expanding the range of cancer types, paving the way for the development of large-scale, general-purpose biological foundation models.
Impact Statement
This work proposes a pretraining framework based on Supervised Graph Contrastive Learning (SupGCL) that learns representations of patient-derived gene regulatory networks (GRNs) using experimentally measured perturbations from gene knockdown experiments as supervision. By aligning synthetic graph perturbations with biologically meaningful perturbations, the proposed approach can better elucidate disease subtype structure in patient GRNs for breast, lung, and colorectal cancers, and is expected to improve performance on multiple downstream tasks such as gene-level functional inference and patient-level prediction. This may facilitate hypothesis generation in cancer research, deepen mechanistic understanding, and improve prioritization of candidate genes and pathways, thereby supporting exploration from basic research to translational studies. All data used in this study are obtained from public resources and do not involve additional collection of personal information. However, the learned representations are statistical in nature and do not directly establish causal effects or guarantee clinical utility.
@@ Acknowledgments
This work was supported by JST Moonshot R&D (JPMJMS2021, JPMJMS2024), JST Research and Development Program for Next-generation Edge AI Semiconductors (JPMJES2511), JSPS KAKENHI (JP25K00148, JP25H02626), and a project (JPNP14004) commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
This work used computational resources of the supercomputer Fugaku provided by RIKEN through the HPCI System Research Project (Project IDs: hp150272, ra000018).
@@
Appendix A Detail Proof of Theorem 4.1
@@ Appendix B Algorithm
Our algorithm follows a standard training loop, consisting of the calculation of $\text{Loss}_{\rm SupGCL}$ and the optimization of $f_{\phi}$ using AdamW. Furthermore, to reduce computational costs, we employ sampling-based estimation of the normalization constant for the calculation of softmax functions $p_{\phi}(b|a)$ and $q_{\phi}(b|a)$, and use importance sampling for the calculation of $\text{Loss}_{\rm node}^{a,b}$ and $\text{Loss}_{\rm Aug}$. For computational performance reasons, we adopted an Augmentation sampling size of 8.
We discuss the computational order and sampling of SupGCL in Appendix F.3.2 and Appendix F.4.2.
@@
Input: Graph Neural Network $f_{\phi}$; patient GRNs $\mathcal{G}_{\rm all}=\{\mathcal{G}^{(i)}\}_{i=1}^{N}$; teacher GRNs $\mathcal{H}_{\rm all}=\{\mathcal{H}_{a}^{(i)}\}_{i\in\mathcal{I}_{a},a\in\mathcal{K}}$
for each $\mathcal{G}\subset\mathcal{G}_{\rm all}$ do
  Sample $a,b\sim\mathrm{U}_{\mathcal{K}}$
  Obtain augmented graphs $\mathcal{G}_{a},\mathcal{G}_{b}$ from $\mathcal{G}$
  Select teacher graphs $\mathcal{H}_{a},\mathcal{H}_{b}$ from $\mathcal{H}_{\rm all}$
  ${\bm{Z}}^{a},{\bm{Z}}^{b}\leftarrow f_{\phi}(\mathcal{G}_{a}),f_{\phi}(\mathcal{G}_{b})$ // Embed target GRNs
  ${\bm{Y}}^{a},{\bm{Y}}^{b}\leftarrow f_{\phi}(\mathcal{H}_{a}),f_{\phi}(\mathcal{H}_{b})$ // Embed teacher GRNs
  $q_{\phi}(b|a)\leftarrow\text{softmax}\!\left(\Big[\frac{\langle{\bm{Z}}^{a},{\bm{Z}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ // Target augmentation distribution
  $p_{\phi}(b|a)\leftarrow\text{softmax}\!\left(\Big[\frac{\langle{\bm{Y}}^{a},{\bm{Y}}^{*}\rangle_{F}}{\tau_{\rm a}}\Big]\right)[b]$ // Teacher augmentation distribution
  $\text{Loss}_{\rm Aug}\leftarrow|\mathcal{K}|\,p_{\phi}(b|a)\big(\log p_{\phi}(b|a)-\log q_{\phi}(b|a)\big)$ // Importance-sampled augmentation loss
  $q_{\phi}(j|i,a,b)\leftarrow\text{softmax}\!\left(\Big[\frac{\langle{\bm{z}}^{a}_{i},{\bm{z}}^{b}_{*}\rangle}{\tau_{n}}\Big]\right)[j]$ // Node-level target distribution
  $\text{Loss}_{\rm Node}^{a,b}\leftarrow\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\log q_{\phi}(i|i,a,b)$
  $\text{Loss}_{\rm SupGCL}\leftarrow|\mathcal{K}|\,p_{\phi}(b|a)\,\text{Loss}_{\rm Node}^{a,b}+\text{Loss}_{\rm Aug}$ // Importance-sampled SupGCL loss
  Update $f_{\phi}$ using AdamW optimizer
end for
Output: Trained Graph Neural Network $f_{\phi}$